# Byzantine-resilient Federated Learning With Adaptivity to Data Heterogeneity

### v3

@@ Abstract
This paper deals with federated learning (FL) in the presence of malicious Byzantine attacks and data heterogeneity. A novel Robust Average Gradient Algorithm (RAGA) is proposed, which leverages the geometric median for aggregation and can freely select the round number for local updating. Different from most existing resilient approaches, which perform convergence analysis based on strongly-convex loss function or homogeneously distributed dataset, we conduct convergence analysis for not only strongly-convex but also non-convex loss function over heterogeneous dataset. According to our theoretical analysis, as long as the fraction of dataset from malicious users is less than half, RAGA can achieve convergence at rate $\mathcal{O}({1}/{T^{2/3-\delta}})$ where $T$ is the iteration number and $\delta\in(0,2/3)$ for non-convex loss function, and at linear rate for strongly-convex loss function. Moreover, stationary point or global optimal solution is proved to obtainable as data heterogeneity vanishes. Experimental results corroborate the robustness of RAGA to Byzantine attacks and verifies the advantage of RAGA over baselines on convergence performance under various intensity of Byzantine attacks, for heterogeneous dataset.
Index Terms:
Federated learning, Byzantine attack, data heterogeneity, robust aggregation, convergence analysis.
@@ I Introduction
The rapid growth of intelligent applications in recent decade promotes a broad adoption of them on various devices or equipments, such as mobile phones, wearable devices, and autonomous vehicles, etc [1, 2, 3]. To enable one intelligent application, raw data from various subscribed users has to be trained together to generate a uniform model. From the perspective of subscribed users, who may distribute broadly, it has a risk of privacy leakage if they offload their raw data to a central server for training [4, 5]. Federated learning (FL) is a distributed training framework and has emerged as a promising solution for this dilemma [6, 7]. In FL, model training is completed by exchanging something about model parameters, such as the model parameter itself or the gradient of loss function with model parameters, between each involving user and one central server iteratively. In each round of iteration, the central server aggregates the model parameters from every user and subsequently broadcasts the aggregated one to all the users. After receiving the broadcasted model parameter, each user performs the role of local updating based on its own data. In such a procedure, there is no exchange of raw data between any user and the central server [8, 9].
However, distributed training like FL has to face robust issue because of the involvement of multiple users. To be specific, due to data corruption, device malfunctioning, or malicious attacks at some users, the information about the model parameter to be uploaded to the central server by these abnormal users may deviate from the expected one [10, 11]. In such a case, the group of abnormal users are called as Byzantine users, and the action of Byzantine users is referred to as Byzantine attacks [12, 13]. As a comparison, the group of normal users are called as Honest users. For a Byzantine attack, it is general to assume the uncertainty of the identity and the population of Byzantine users. What is more, the attack initiated by Byzantine users could be arbitrarily malicious [14, 15]. With such a setup, the training performance of FL will be surely degraded and aggregation strategies resistant to Byzantine attacks have been concerned in literature [16].
On the other hand, the training process of FL also faces the challenges of data heterogeneity among multiple users. This heterogeneity comes from the diverse distribution of local data at multiple users, since the local data may be collected by each FL user under its specific environment, habit, or preference [17]. In such a case, data distributions over multiple users are always assumed to be non independently and identically distributed (non-IID). The presence of data heterogeneity has already been demonstrated to cause local drift of user model, and thus, degradation of the global model’s performance. How to be adaptive to data heterogeneity is also an important issue and has been studied extensively in the area of FL [18].
@@ I-A Related Works
In related works, aggregation strategies are designed to be robust to the Byzantine attack under various setup of system assumptions, whose performance is then justified by convergence analysis and/or experimental results. The system assumptions usually relate to the nature of the data (IID or non-IID), the nature of the loss function (strongly-convex or non-convex), and the way of local updating (one-step or multiple steps), etc. For these assumptions, the data heterogeneity and non-convex loss function make FL framework more general, and multiple steps of local updating waives each user from frequent communication with the central server. The Byzantine-resilient aggregation strategies in the literature of FL under diverse system assumptions have been summarized in Table I.
@@
References Algorithm Function Type Data Heterogeneity Local Updating Optimality Gap Other Assumption
[16] median - IID one step - -
[19] trimmed mean non-convex strongly-convex IID one step non-zero Bounded gradient
[20] iterative filtering strongly-convex IID one step non-zero -
[21] RANGE non-convex strongly-convex IID one step non-zero -
[22] - non-convex strongly-convex IID little data heterogeneity multiple step non-zero Sub-exponential distribution
[23] - strongly-convex IID multiple step non-zero Sub-exponential distribution
[24] Krum non-convex IID one step non-zero -
[25] FLOA non-convex IID one step zero Bounded gradient Unbiased gradient
[26] RSA strongly-convex non-IID one step non-zero Bounded gradient
[27] RFA strongly-convex non-IID multiple steps non-zero -
[28] Byrd-SAGA strongly-convex IID one step non-zero Bounded gradient Unbiased gradient
[29] BROADCAST strongly-convex IID one step non-zero Bounded gradient Unbiased gradient
This work RAGA non-convex strongly-convex non-IID multiple steps non-zero Bounded gradient
@@
As of the aggregation strategies in related literature, by leveraging median [16], trimmed mean [19], or iterative filtering [20], the associated FL system can tolerate a small fraction of Byzantine users. To be more robust to Byzantine attacks, some methods such as Robust Aggregating Normalized Gradient Method (RANGE) [21], a Huber Loss Minimization [22], semi verified mean estimation [23], Krum [24], Federated Learning Over the Air (FLOA) [25], and Robust Stochastic Aggregation (RSA) [26], have been proposed. Specifically, RANGE takes the coordinate-wise median of uploaded stochastic gradients as the aggregated one. [22] improve the trimmed mean on non-IID datasets by utilizing a multi-dimensional Huber loss function to aggregate uploaded messages. [23] leverges semi verified mean estimation to filter out the messages from Byzantine attackers before aggregation. Krum selects a stochastic gradient as the global one, which has the shortest cumulative distance from a group of stochastic gradients that are most closely distributed, FLOA requires each user to normalize the associated local gradient to be a vector with zero mean and unit variance before offloading it to the central server, RSA penalizes the difference between local model parameters and global model parameters to isolate Byzantine users.
@@
Although the above aggregation strategies can be resistant to Byzantine attacks at some extent, they may still fail to work in front of heavy Byzantine attacks without identifying the Byzantine attacker: The attacks with the fraction of Byzantine users being large or the amplitude of uploaded vectors by Byzantine users being high. Recently, geometric median appears as a good candidate for aggregation because of its ability to naturally tolerate a larger fraction of strong Byzantine attacks [30]. Based on geometric median, some algorithms such as Robust Federated Aggregation (RFA) [27], Byzantine attack resilient distributed Stochastic Average Gradient Algorithm (Byrd-SAGA) [28], and Byzantine-RObust Aggregation with gradient Difference Compression And STochastic variance reduction (BROADCAST) [29], have been proposed. To be specific, geometric median is leveraged to aggregate the uploaded vectors for RFA, Byrd-SAGA, and BROADCAST. Differently, each user in RFA selects the trimmed mean of the model parameters over multiple local updating as the uploaded vector, while Byrd-SAGA and BROADCAST utilize the SAGA [31] method to generate the vectors for aggregation first and then upload them in a lossless or compressive way.
@@
In aforementioned Byzantine-resilient aggregation strategies, there is little work claimed to be adaptive to data heterogeneity, which however is a non-negligible issue. RFA in [27] is an exception. However, RFA only investigates strongly-convex loss function. With such a convexity setup, RFA shows that it is can converge to a neighborhood of the optimal solution only if the number of local updates for each involving user is larger than then a threshold so as to fulfill a necessary inequality. However, this restricts the flexibility of selecting the number of local updates and maybe computationally intensive when these users have limited computation capability.
@@ I-B Motivations and Contributions
In this paper we investigate Byzantine-resilient FL with adaptivity to data heterogeneity, which was rarely studied in literature. Different from RFA, which is the unique work dealing with both Byzantine attack and data heterogeneity in FL by now but only considers strongly-convex loss function, we investigate not only strongly-convex but also non-convex loss function. Moreover, with regard to strongly-convex loss function, not like RFA that requires the step number of local updates to be above some threshold (which is a prerequisite for promising convergence in RFA), we assume arbitrary selection of it, which is more general. With our featured system setup, we propose a novel FL structure, RAGA. Specifically, the RAGA leverages the geometric median to aggregate the local gradients offloaded by the users after multiple rounds of local updating. Associated analysis of convergence and robustness are then followed for both non-convex and strongly-convex loss functions. Our main contributions are summarized as follows:
  • Algorithmically, we propose a new FL aggregation strategy, named RAGA, which has the flexibility of arbitrary selection of local update’s step number, and is robust to both Byzantine attack and data heterogeneity.
  • Theoretically, we established a convergence guarantee for our proposed RAGA under not only non-convex but also strongly-convex loss function. Through rigorous proof, we show that the established convergence can be promised as long as the fraction of dataset from Byzantine users is less than half. The achievable convergence rate can be at $\mathcal{O}({1}/{T^{2/3-\delta}})$ where $T$ is the iteration number and $\delta\in(0,2/3)$ for non-convex loss function, and at linear for strongly-convex loss function. Moreover, stationary point and global optimal solution are shown to be obtainable as data heterogeneity vanishes, for non-convex and strongly-convex loss function, respectively.
  • Empirically, we conduct extensive experiments to evaluate the performance of our proposed RAGA. With non-IID dataset, our proposed RAGA is verified to be robust to various Byzantine attack level. Moreover, compared with baseline methods, our proposed RAGA can achieve higher test accuracy and lower training loss.
@@ II Problem Statement
$F(\bm{w})\triangleq\frac{1}{\sum_{m\in\mathcal{M}}S_{m}}\sum_{m\in\mathcal{M}}\sum_{\bm{s}\in\mathcal{S}_{m}}f(\bm{w},\bm{s})$ (1)
@@
For the global loss function $F(\bm{w})$ and local loss function $F_{m}(\bm{w})$ for $m\in\mathcal{M}$, we adopt assumptions from the following ones.
Assumption 1 (Lipschitz Continuity)
$\displaystyle f(\bm{w}_{1},\bm{s})-f(\bm{w}_{2},\bm{s})\leqslant\left\langle\nabla f(\bm{w}_{2},\bm{s}),\bm{w}_{1}-\bm{w}_{2}\right\rangle+\frac{L}{2}\left\lVert\bm{w}_{1}-\bm{w}_{2}\right\rVert^{2},$ (4)
which is equivalent to the following inequality,
With (4) or (5), it can be derived that the global loss function $F(\bm{w})$ and local loss function $F_{m}(\bm{w}),\forall m\in\mathcal{M}$ both have $L$-Lipschitz continuity.
Assumption 2 (Strongly-Convex)
Similarly, with (6), $F(\bm{w})$ and $F_{m}(\bm{w}),\forall m\in\mathcal{M}$ can be derived to be both $\mu$ strongly-convex.
Assumption 3 (Local Unbiased Gradient)
For $\forall m\in\mathcal{M}$, suppose a subset of dataset $\mathcal{S}_{m}$ is selected randomly, which is denote as $\xi_{m}$, define
the unbiased gradient [32, 33] implies that
$\mathbb{E}\left\{\nabla F_{m}(\bm{w};\xi_{m})\right\}=\nabla F_{m}(\bm{w}),\forall m\in\mathcal{M}.$ (8)
Assumption 4 (Bounded Inner Variance)
For every user $m$ and $\bm{w}\in\mathbb{R}^{p}$, the variance of stochastic local gradient $\nabla F_{m}(\bm{w};\xi_{m})$ is upper-bounded [32, 33] by
$\displaystyle\text{Var}(\nabla F_{m}(\bm{w};\xi_{m}))\leqslant\sigma^{2},\forall m\in\mathcal{M}.$ (9)
Assumption 5 (Bounded Gradient)
For every user $m$ and $\bm{w}\in\mathbb{R}^{p}$, the ideal local gradient $\nabla F_{m}(\bm{w})$ is upper-bounded [32, 33] by
$\displaystyle\left\lVert\nabla F_{m}(\bm{w})\right\rVert\leqslant G,\forall m\in\mathcal{M}.$ (10)
@@
Step 1 (Local Update): For $m$th user, it calculates the local gradient parameter $\nabla F_{m}(\bm{w}^{t})$ from its local dateset $\mathcal{S}_{m}$ and global model parameter $\bm{w}^{t}$.
@@ III Proposed Aggregation Strategy: Robustness Average Gradient Algorithm (RAGA)
$\bm{z}_{m}^{t}=\left\{\begin{aligned} &{\bm{g}_{m}^{t},}&&{m\in\mathcal{M}\setminus\mathcal{B}}\\ &{\bm{\star},}&&{m\in\mathcal{B}}\end{aligned}\right.$ (12)
@@
$\displaystyle\bm{w}_{m}^{t,k}=\bm{w}_{m}^{t,k-1}-\eta_{m}^{t,k}\cdot\nabla F_{m}(\bm{w}_{m}^{t,k-1};\xi_{m}^{t,k}),k=1,2,\dots,K^{t}$ (13)
@@
$\displaystyle\bm{w}^{t+1}$ $\displaystyle=\bm{w}^{t}-\eta^{t}\cdot\bm{z}^{t}$ (14)
@@
$\displaystyle{\rm geomed}\left(\{\bm{z}_{m}^{t}|m\in\mathcal{M}\}\right)\triangleq\bm{z}^{t,*}=\arg\min_{y}\sum_{i=1}^{M}\frac{S_{i}}{\sum_{j=1}^{M}S_{j}}\left\lVert y-\bm{z}_{i}^{t}\right\rVert.$ (15)
The problem in (15) is actually a convex optimization problem and an $\epsilon$-optimal solution is obtainable by resorting to Weiszfeld algorithm [34]. Hence the $\bm{z}^{t}$ fulfills the following condition
@@
and is also denoted as ${\rm geomed}\left(\{\bm{z}_{m}^{t}|m\in\mathcal{M}\},\epsilon\right)$ in the sequel. Then the central server broadcasts the global model parameter $\bm{w}_{t+1}$ to all the users for preparing the calculation in $t+1$ round of iteration.
@@
1 The central server initializes the model parameter $\bm{w}^{1}$ and sends it to all the users.

### v4

# Federated Learning Resilient to Byzantine Attacks and Data Heterogeneity

@@ Abstract
This paper addresses federated learning (FL) in the context of malicious Byzantine attacks and data heterogeneity. We introduce a novel Robust Average Gradient Algorithm (RAGA), which uses the geometric median for aggregation and allows flexible round number for local updates. Unlike most existing resilient approaches, which base their convergence analysis on strongly-convex loss functions or homogeneously distributed datasets, this work conducts convergence analysis for both strongly-convex and non-convex loss functions over heterogeneous datasets. The theoretical analysis indicates that as long as the fraction of the data from malicious users is less than half, RAGA can achieve convergence at a rate of $\mathcal{O}({1}/{T^{2/3-\delta}})$ for non-convex loss functions, where $T$ is the iteration number and $\delta\in(0,2/3)$. For strongly-convex loss functions, the convergence rate is linear. Furthermore, the stationary point or global optimal solution is shown to be attainable as data heterogeneity diminishes. Experimental results validate the robustness of RAGA against Byzantine attacks and demonstrate its superior convergence performance compared to baselines under varying intensities of Byzantine attacks on heterogeneous datasets.
@@ I Introduction
The rapid growth of intelligent applications in recent decade promotes a broad adoption of them on various devices or equipments, such as mobile phones, wearable devices, and autonomous vehicles, etc [1, 2, 3]. To enable one intelligent application, raw data from various subscribed users has to be trained together to generate a uniform model. From the perspective of subscribed users, who may distribute broadly, it has a risk of privacy leakage if they offload their raw data to a central server for training [4, 5]. Federated learning (FL) is a distributed training framework and has emerged as a promising solution for this dilemma [6]. In FL, model training is completed by exchanging something about model parameters, such as the model parameter itself or the gradient of the loss function with respect to model parameters, between each involving user and one central server iteratively. In each round of iteration, the central server aggregates the model parameters from every user and subsequently broadcasts the aggregated one to all the users. After receiving the broadcasted model parameter, each user performs the role of local updating based on its own data. In such a procedure, there is no exchange of raw data between any user and the central server [7, 8].
However, distributed training paradigms such as FL inherently encounter robustness challenges stemming from the participation of multiple heterogeneous users. To be specific, due to data corruption, device malfunctioning, or malicious attacks at some users, the information about the model parameter to be uploaded to the central server by these abnormal users may deviate from the expected one [9, 10]. In such a case, the group of abnormal users are called as Byzantine users, and the action of Byzantine users is referred to as Byzantine attacks [11, 12]. As a comparison, the group of normal users are called as Honest users. For a Byzantine attack, it is general to assume the uncertainty of the identity and the population of Byzantine users. What is more, the attack initiated by Byzantine users could be arbitrarily malicious [13, 14]. With such a setup, the training performance of FL will be surely degraded and aggregation strategies resistant to Byzantine attacks have been concerned in literature [15].
On the other hand, the training process of FL faces challenges of data heterogeneity among users. This heterogeneity comes from the diverse distribution of local data, as the local data may be collected by each FL user in their specific environmental conditions, habits, or preferences [16]. In such a case, data distributions over multiple users are always assumed to be non independently and identically distributed (non-IID). The presence of data heterogeneity has already been demonstrated to cause local drift of user model, and thus, degradation of the global model’s performance. How to be adaptive to data heterogeneity is also an important issue and has been studied extensively in the area of FL [17].
@@ I-A Related Works
In the field of FL, numerous studies have considered a wide range of factors. However, research on aggregation strategies that are robust to Byzantine attacks in association with rigorous convergence analysis remains limited, primarily due to the significant challenges such attack introduces into the aggregation process. The system assumptions, which serve as an essential precondition for convergence analysis, usually relate to the nature of the data (IID or non-IID), the nature of the loss function (strongly-convex or non-convex), and the way of local updating (one-step or multiple steps), etc. For these assumptions, the data heterogeneity and non-convex loss function make FL framework more general, and multiple steps of local updating waives each user from frequent communication with the central server. The Byzantine-resilient aggregation strategies in the literature of FL under diverse system assumptions have been summarized in Table I.
@@
References Algorithm Function Type Data Heterogeneity Local Updating Optimality Gap
[15] median - IID one step -
[18] trimmed mean
non-convex
strongly-convex
IID one step non-zero
[19] RANGE
non-convex
strongly-convex
IID one step non-zero
[20] -
non-convex
strongly-convex
little data heterogeneity one step non-zero
[21] iterative filtering strongly-convex IID one step non-zero
[22] - strongly-convex IID one step non-zero
[23] Krum non-convex IID one step non-zero
[24] BEV-SGD non-convex IID one step zero
[25] RSA strongly-convex non-IID one step non-zero
[26] MCA strongly-convex non-IID one step non-zero
[27] RFA strongly-convex non-IID multiple steps non-zero
[28] Byrd-SAGA strongly-convex IID one step non-zero
[29] BROADCAST strongly-convex IID one step non-zero
[30] SDEA - non-IID multiple steps -
[31] RAGE
non-convex
strongly-convex
non-IID multiple steps non-zero
This work RAGA
non-convex
strongly-convex
non-IID multiple steps non-zero
@@
As of the aggregation strategies in related literature, by leveraging median [15] or iterative filtering [21], the associated FL system can tolerate a small fraction of Byzantine users. To be more robust to Byzantine attacks, some methods such as trimmed mean [18], Robust Aggregating Normalized Gradient Method (RANGE) [19], a Huber Loss Minimization [20], semi verified mean estimation [22], Krum [23], Best Effort Voting SGD (BEV-SGD) [24], Robust Stochastic Aggregation (RSA) [25], Maximum Correntropy Aggregation (MCA) [26], Self-Driven Entropy Aggregation (SDEA) [30], and Robust Accumulated Gradient Estimation (RAGE) [31] have been proposed. Specifically, [18] aggregates the upload vectors by calculating their coordinate-wise median or coordinate-wise trimmed mean based on [15]. RANGE takes the median of uploaded vectors from each client as the aggregated one, and each uploaded vector is supposed to be the median of the previously transmitted gradients. [20] improve the trimmed mean on non-IID datasets by utilizing a multi-dimensional Huber loss function to aggregate uploaded vectors. [22] utilizes semi-verified mean estimation to filter out vectors from Byzantine attackers before aggregation, building on the iterative filtering method. Krum [23] selects a stochastic gradient as the global one, knowing the number of Byzantine attackers, by choosing the gradient with the shortest cumulative distance from a group of the most closely distributed stochastic gradients, BEV-SGD [24] requires each user to normalize the associated local gradient to be a vector with zero mean and unit variance before offloading it to the central server, RSA [25] penalizes the difference between local model parameters and global model parameters to isolate Byzantine users, while MCA [26] utilizes its the maximum correntropy of the model parameters’ distribution characteristics and compute the central value to update the global model parameter. Moreover, the SDEA method [30] partially leverages a public dataset to enhance global model training and mitigate Byzantine attacks, with its effectiveness validated through empirical experiments. In contrast, the RAGE approach [31] enables multiple local gradient updates under heterogeneous data distributions. While theoretically well-founded, this method assumes that the Byzantine ratio, i.e., the proportion of malicious participants, remains below one-third.
@@
Although the above aggregation strategies can be somewhat resistant to Byzantine attacks at some extent, some of them may still fail to work in front of heavy Byzantine attacks [32, 33]. Recently, the geometric median has emerged as a promising candidate for aggregation due to its ability to naturally tolerate a larger fraction of strong Byzantine attacks without any additional conditions[34]. Based on geometric median, some algorithms such as Robust Federated Aggregation (RFA) [27], Byzantine attack resilient distributed Stochastic Average Gradient Algorithm (Byrd-SAGA) [28], and Byzantine-RObust Aggregation with gradient Difference Compression And STochastic variance reduction (BROADCAST) [29], have been proposed. To be specific, geometric median is leveraged to aggregate the uploaded vectors for RFA, Byrd-SAGA, and BROADCAST. Differently, each user in RFA selects the trimmed mean of the model parameters over multiple local updating as the uploaded vector, while Byrd-SAGA and BROADCAST utilize the SAGA [35] method to generate the vectors for aggregation first and then upload them in a lossless or compressive way.
@@
In the aforementioned geometric median aggregation strategies, there has been little work claiming to be adaptive to data heterogeneity, though this remains a non-negligible issue. RFA in [27] is an exception. However, RFA only investigates strongly-convex loss function. With such a convexity setup, RFA shows that it is can converge to a neighborhood of the optimal solution only if the number of local updates for each involving user is larger than then a threshold so as to fulfill a necessary inequality. However, this restricts the flexibility of selecting the number of local updates and maybe computationally intensive when these users have limited computation capability.
@@ I-B Motivations and Contributions
In this paper we investigate Byzantine-resilient FL with adaptivity to data heterogeneity, which was only studied in [25], [26], [27], [30], and [31], referred to as RSA, MCA, RFA, SDEA, and RAGE, respectively. Different from RSA, MCA, and RFA, which merely considers strongly-convex loss function, we investigate not only strongly-convex but also non-convex loss function. Specifically, in departure from RFA, which also uses geometric median for aggregation but has a convergence condition that necessitates local update steps exceeding a theoretically established threshold, our architecture introduces a flexible round number of local updates scheme, thereby eliminating this fundamental constraint while maintaining convergence guarantees.
While SDEA is effective in practice, it lacks a formal theoretical convergence analysis and depends on the availability of public data, which may not always be practical or feasible. In contrast to SDEA, our approach does not assume access to any public data and provides rigorous theoretical guarantees. Moreover, unlike RAGE, which requires the Byzantine ratio to be no greater than one-third, our method can tolerate a higher Byzantine ratio, potentially approaching one-half. With our featured system setup, we propose a novel FL structure, RAGA. Specifically, the RAGA leverages the geometric median to aggregate the local gradients offloaded by the users after multiple rounds of local updating. Associated analysis of convergence and robustness are then followed for both non-convex and strongly-convex loss functions. Our main contributions are summarized as follows:
  • Algorithmically, we propose a new FL aggregation strategy, named RAGA, which has the flexibility of arbitrary selection of local update’s step number, and is robust to both Byzantine attack and data heterogeneity, and does not rely on the use of a public dataset.
  • Theoretically, we establish a convergence guarantee for our proposed RAGA under not only non-convex but also strongly-convex loss function. Through rigorous proof, we show that the established convergence can be promised as long as the fraction of dataset from Byzantine users is less than half. The achievable convergence rate can be at $\mathcal{O}({1}/{T^{2/3-\delta}})$ where $T$ is the iteration number and $\delta\in(0,2/3)$ for non-convex loss function, and at linear for strongly-convex loss function. Moreover, stationary point and global optimal solution are shown to be obtainable as data heterogeneity, algorithmic error, and stochastic local gradient variance diminish, for non-convex and strongly-convex loss function, respectively.
  • Empirically, we conduct extensive experiments to evaluate the performance of our proposed RAGA. With non-IID dataset, our proposed RAGA is verified to be robust to various Byzantine attack types and ratios. Moreover, compared with baseline methods, our proposed RAGA can achieve higher test accuracy.
@@ II Problem Statement
$F(\bm{w})\triangleq\frac{1}{\sum_{m\in\mathcal{M}}S_{m}}\sum_{m\in\mathcal{M}}\sum_{\bm{s}\in\mathcal{S}_{m}}f(\bm{w},\bm{s}),$ (1)
@@
Step 1 (Local Update): For $m$th user, it calculates the local gradient parameter $\nabla F_{m}(\bm{w}^{t})$ from its local dataset $\mathcal{S}_{m}$ and global model parameter $\bm{w}^{t}$.
@@ III Proposed Aggregation Strategy: Robustness Average Gradient Algorithm (RAGA)
$\bm{z}_{m}^{t}=\left\{\begin{aligned} &{\bm{g}_{m}^{t},}&&{m\in\mathcal{M}\setminus\mathcal{B}}\\ &{\bm{\star},}&&{m\in\mathcal{B}}\end{aligned}\right$., (5)
@@
$\displaystyle\bm{w}_{m}^{t,k}=\bm{w}_{m}^{t,k-1}-\eta_{m}^{t,k}\cdot\nabla F_{m}(\bm{w}_{m}^{t,k-1};\xi_{m}^{t,k}),$
$\displaystyle k=1,2,\dots,K^{t},$ (6)
@@
$\displaystyle\bm{w}^{t+1}$ $\displaystyle=\bm{w}^{t}-\eta^{t}\cdot\bm{z}^{t},$ (7)
@@
$\displaystyle\bm{z}^{t,*}$ $\displaystyle\triangleq{\rm geomed}\left(\{\bm{z}_{m}^{t}|m\in\mathcal{M}\}\right)$
$\displaystyle=\arg\min_{y}\sum_{i=1}^{M}\frac{S_{i}}{\sum_{j=1}^{M}S_{j}}\left\lVert y-\bm{z}_{i}^{t}\right\rVert.$ (8)
The problem in (III) is actually a convex optimization problem and an $\epsilon$-optimal solution is obtainable by resorting to Weiszfeld algorithm [36]. Hence the $\bm{z}^{t}$ fulfills the following condition
@@
and is also denoted as ${\rm geomed}\left(\{\bm{z}_{m}^{t}|m\in\mathcal{M}\},\epsilon\right)$ in the sequel. Then the central server broadcasts the global model parameter $\bm{w}^{t+1}$ to all the users for preparing the calculation in $t+1$ round of iteration.
@@
1The central server initializes the model parameter $\bm{w}^{1}$ and sends it to all the users.